OpenStudy (anonymous):

Identify the x-intercepts in the graph y = 3x^2 + 19x – 40.

OpenStudy (vishweshshrimali5):

y = 0 thus, 3x^2 + 19x - 40 = 0 => x = 5/3, x = -8 thus the points are (5/3, 0) and (-8,0)

OpenStudy (anonymous):

You're getting a lot of medals today lol. Thanks again <3